OpenStudy (anonymous):

On a school map, the points (8, -4), (8, -10), (2, -4), and (2, -10) represent the four corners of a sand park.

OpenStudy (anonymous):

What are the coordinates of the center of the sand park?

OpenStudy (anonymous):

all you have to do is drwaw the diagram and find the mid point when you compare two sides..the formular is (x,y)=(x1+x2)/2 , (y1+y2)/2 i used (2,-4) and (8,-10) the center of the sand park (x,y)=(2+8)/2 , (-4+(-10)) x,y=10/2 , -14/2 x,y=5,-7 therefore the centre of the sand is (5,-7)...dont mind my diagram